OpenStudy (anonymous):

two friends Aman and Raman are going to Ahmedabad from Delhi via Jaipur. Aman started at 8 am from Delhi and maintained a speed of 50 kmph upto Jaipur and 60 km in between Jaipur and Ahmedabad whereas Raman left Delhi at 9.45 am and maintained an average speed of 60 kmph in between Delhi and Jaipur and 80 kmph between Jaipur and Ahmedabad. Both Aman and Raman reached Ahmedabad at 4 pm. What is the distance between Jaipur and Ahmedabad?

OpenStudy (anonymous):

total time taken by aman = 8 hours distance between D to J = x J to A= y 8 = 50x+60y now by the same method form the equations for raman

OpenStudy (anonymous):

it will become a system of equation in two variables

OpenStudy (anonymous):

6.25=60x+80y-2nd x=67/160 y=137/400 x+y=609/800

OpenStudy (anonymous):

is it right

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

its not the answer.i think becoz he maintained inbetween avg speed

OpenStudy (tkhunny):

There is enough information. I did it this way, using Distance = Rate * Time four different times. DJ - Distance from Delhi to Jaipur t - Aman's time for DJ s - Raman's time for DJ JA - Distance from Jaipur to Ahmedabad Aman - Total Trip, 8 Hours DJ = t*50 JA = (8-t)*60 Raman - Total Trip, 6.25 hours DJ = s*60 JA = (6.25-s)*80 Four equations. Four unknown quantities. We're almost done.